OpenStudy (anonymous):

Which ratio correctly compares 4 feet to 36 inches? A. 1\10 B. 1\9 C. 3\4 D. 4\3

OpenStudy (katherinkoon):

i think it's b

rebeccaxhawaii (rebeccaxhawaii):

4/36=1/9

OpenStudy (anonymous):

thanksss

OpenStudy (anonymous):

1 feet = 12 inches, 4 feet= 48 inches 4feet/36 inches= 48/36=4/3=1.33

rebeccaxhawaii (rebeccaxhawaii):

no problem

OpenStudy (anonymous):

D 4/3 :) Convert both to the same measure, and you have your answer. 4 feet is 48 inches, so it could be 48:36. But 36 inches is also 3 ft, so a simpler ratio would be 4:3. 4:3 is the same as 4/3.